OpenStudy (anonymous):

how do you write an equation, such as "y+6 = (x+4)?"

OpenStudy (anonymous):

y=x+4-6 y=x-2 or x-y=2

OpenStudy (anonymous):

y=x-2

OpenStudy (anonymous):

how did u get that?

OpenStudy (anonymous):

Subtract "6" from the right side of the given equation sooo y+6 = (x+4) ----> y = (x+4) - 6 ----> y = x-2 Understand?